OpenStudy (anonymous):

Silver scooter finds that it costs $300 to produce each scooter and the fixed costs are $750. The price is given by p=900-x, where p is the price in dollars at which exactly x scooters will be sold. Find the quantity of scooters that the company should produce and the price it should charge to maximize profit. Find the maximum profit.

OpenStudy (anonymous):

the cost to produce x scooters is C = 750 + 300x Price of one scooter is P = 900-x profit = sales price of x scooters - cost profit = x(900-x) - 750 - 300x profit = 900x - x^2 - 750 - 300x profit = -x^2 +600x -750 maximum profit is got by differentiating the above equation with respect to x and setting it to 0 d(profit) / dx = -2x +600 -2x+600 = 0 2x = 600 x = 300 so the company must produce 300 scooters to maximize profit. the price would then be 900 - x = 900 - 300 = $600 the maximum profit is 300(900-300) - 750 - 300*300

OpenStudy (anonymous):

= $89250